Consider a portfolio made up of two stocks, S1 an S2. The variance of the 1-year returns on the two stocks are denoted s12 and s22

 

[a] Assume we construct a portfolio consisting of long positions in the two stocks. Explain why the variance of the 1-year return on the portfolio can’t exceed the larger of s12 and s22.

 

Suppose we invest $150 in a stock S1:

$100 of our own capital

$50 of borrowed 1-year money (at an interest rate of 3%)

 

Let R1 denote the 1-year return on S1. Let E[R1] = m. 

 

[b] In terms of s12 what’s the expected value and variance of the 1-year return on our $100 investment?


1
Expert's answer
2021-05-07T12:25:20-0400

 The variance of the 1-year returns on the two stocks are denoted s12 and s22

 

(a) When we construct a portfolio consisting of long positions in the two stocks. variance of the 1-year return on the portfolio can't exceed because the portfolio variance is the maximum variance occur on a investment.


Investment I=$150

"R_1" return after 1 year on "s_1"


(b)  expected value "= E[R_1]-\sqrt{s_1^2}"


and variance of the 1-year return "= s_1^2-\dfrac{3\times 50}{100}=s_1^2-1.5"